OpenStudy (anonymous):

Solve the inequality x+7≤ 34 and describe the graph of the solutions. Show all work.

HanAkoSolo (jamierox4ev3r):

to solve this inequality, just solve for one variable. basically, isolate x all by itself. treat the inequality like an equal sign for now. Does that makes sense? @MeMe22 ?

HanAkoSolo (jamierox4ev3r):

I mean, does it make sense thus far? :P

HanAkoSolo (jamierox4ev3r):

hello? lol

OpenStudy (anonymous):

Sorry I am back and yes it does

HanAkoSolo (jamierox4ev3r):

so what do you get once you fully isolate for x? @MeMe22 ? and how would you graph it? if you still don't know, feel free to ask me :)

OpenStudy (anonymous):

Im not sure

HanAkoSolo (jamierox4ev3r):

so if x+7\(\le\)34, solve like a regular equation: -7 -7 ----------- x\(\le\)27

HanAkoSolo (jamierox4ev3r):

then if this only equals x and it is not in full y=mx+b formula, the line can either be straight or horizontal. Which one do you think it is @MeMe22

HanAkoSolo (jamierox4ev3r):

Hey I have to leave soon, but if you still have questions, feel free to hit me up. My name is Jamie :)